# A Hybrid Mean Field Framework for Aggregators Participating in Wholesale Electricity Markets

### v1

@@ 4.1   Components of An Aggregator’s Problem
To promote exploration and improve learning stability, we incorporate entropy regularization into the reward function. Specifically, the single-period reward is modified as: $r^{\text{REG}}_{n}(s,a,\bm{\lambda}^{n})=r_{n}(s,a,\bm{\lambda}^{n})-\Omega(\pi(\cdot\mid s,\bm{\lambda}^{n})),$ where $\Omega(\cdot)$ is a $\rho$-strongly convex regularization function. A common choice is the negative entropy: $\Omega(\pi(\cdot\mid s,\bm{\lambda}^{n}))=\alpha\sum_{a\in\mathcal{A}}\pi(a\mid s,\bm{\lambda}^{n})\log\pi(a\mid s,\bm{\lambda}^{n}),$ with $\alpha>0$ controlling the strength of regularization. The corresponding regularized value function is: $V^{\text{REG}}_{n}(s^{n}_{0},\pi^{n}_{t},\bm{\lambda}^{n})=\mathbb{E}\left[\sum_{\tau=0}^{\infty}\gamma_{n}^{\tau}r^{\text{REG}}_{n}(s^{n}_{\tau},a^{n}_{\tau},\bm{\lambda}^{n})\right].$ Each aggregator seeks to learn an optimal policy $\pi^{n*}_{t}$ that maximizes $V^{\text{REG}}_{n}$ at each time $t$.
@@ A.1   Uniqueness of Optimal Policy under Regularization
We argue that the expected reward $\mathbb{E}[r_{n}(s,a,\bm{\lambda})]$ is linear w.r.t. $\pi^{n}$. In fact, the linearity is a direct consequence of the Lebesgue measure by viewing the distribution $\pi^{n}$ as the measure function. Then the sum of a linear function and a $\rho$-strongly concave function preserves the $\rho$-strong concavity. Thus, the $\arg\max_{\pi}V^{\text{REG}}_{n}$ admits a unique solution. ∎
@@ A.2   Proof of Theorem 1
Proof.
In different rounds, we solve different RL problems by dealing with different $Q$-functions. Given $\bm{\lambda}$, we define
$Q_{\bm{\lambda}}(s,a)=r^{\text{REG}}(s,a,\bm{\lambda})+\gamma\sum_{s^{\prime}}Q_{\bm{\lambda}}^{\star}(s^{\prime})P(s^{\prime}|s,a),$ (20)
where $Q_{\bm{\lambda}}^{\star}(s)=\max_{a}Q_{\bm{\lambda}}(s,a)$ for all $s$. Then, for any $\bm{\lambda}_{1},\bm{\lambda}_{2}$, we have
@@
$\hskip 10.0pt=\max_{s,a}\Big{|}r^{\text{REG}}(s,a,\bm{\lambda}_{1})-r^{\text{REG}}(s,a,\bm{\lambda}_{2})+\gamma\sum_{s^{\prime}}Q_{\bm{\lambda}_{1}}^{\star}(s^{\prime})P(s^{\prime}|s,a)-\gamma\sum_{s^{\prime}}Q_{\bm{\lambda}_{2}}^{\star}(s^{\prime})P(s^{\prime}|s,a)\Big{|}$
$\hskip 10.0pt\leq\max_{s,a}\Big{\{}|r(s,a,\bm{\lambda}_{1})-r(s,a,\bm{\lambda}_{2})|+\gamma\Big{|}\sum_{s^{\prime}}Q_{\bm{\lambda}_{1}}^{\star}(s^{\prime})P(s^{\prime}|s,a)-\sum_{s^{\prime}}Q_{\bm{\lambda}_{2}}^{\star}(s^{\prime})P(s^{\prime}|s,a)\Big{|}\Big{\}}$
$\hskip 10.0pt\leq L_{r}\|\bm{\lambda}_{1}-\bm{\lambda}_{2}\|_{1}+\gamma\|Q_{\bm{\lambda}_{1}}^{\star}-Q_{\bm{\lambda}_{2}}^{\star}\|_{\infty},$
where the first term follows the continuity of the reward functions and the second term follows the fact that $P$ is probability. Then by rearranging the terms, we get that $\|Q_{\bm{\lambda}_{1}}^{\star}-Q_{\bm{\lambda}_{2}}^{\star}\|_{\infty}\leq\frac{L_{r}}{1-\gamma}\|\bm{\lambda}_{1}-\bm{\lambda}_{2}\|_{1}$. Notice that Equation (20) can be rewritten as follows:
@@
where $q_{\bm{\lambda}}^{s}=r(s,\cdot,\bm{\lambda})+\gamma\sum_{s^{\prime}}Q_{\bm{\lambda}}^{\star}(s^{\prime})P(s^{\prime}|s,\cdot)$ for any $s$, which is $\left(L_{r}+\frac{\gamma L_{r}}{1-\gamma}\right)$-Lipschtiz continuous with respect to $\bm{\lambda}$, as shown below:
@@
$\hskip 10.0pt\leq L_{r}\|\bm{\lambda}_{1}-\bm{\lambda}_{2}\|_{1}+\gamma\max_{a}\Big{|}\sum_{s^{\prime}}Q_{\bm{\lambda}_{1}}^{\star}(s^{\prime})P(s^{\prime}|s,a)-\sum_{s^{\prime}}Q_{\bm{\lambda}_{2}}^{\star}(s^{\prime})P(s^{\prime}|s,a)\Big{|}$
$\hskip 10.0pt\leq L_{r}\|\bm{\lambda}_{1}-\bm{\lambda}_{2}\|_{1}+\gamma\|Q_{\bm{\lambda}_{1}}^{\star}-Q_{\bm{\lambda}_{2}}^{\star}\|_{\infty}=\left(L_{r}+\frac{\gamma L_{r}}{1-\gamma}\right)\|\bm{\lambda}_{1}-\bm{\lambda}_{2}\|_{1}.$
We finally apply the Fenchel conjugate properties from Proposition 3. For any $s\in\mathcal{S}$, we write $\Gamma^{n}_{1}(s,\bm{\lambda})=\nabla\Omega^{\star}(q_{\bm{\lambda}}^{s})$ where $\Omega^{\star}$ is the Fenchel conjugate of $\Omega$. Then,
$\|\Gamma^{n}_{1}(s,\bm{\lambda}_{1})-\Gamma^{n}_{1}(s,\bm{\lambda}_{2})\|_{1}\leq\frac{1}{\rho}\|q_{\bm{\lambda}_{1}}^{s}-q_{\bm{\lambda}_{2}}^{s}\|_{\infty}=\frac{1}{\rho}\left(L_{r}+\frac{\gamma L_{r}}{1-\gamma}\right)\|\bm{\lambda}_{1}-\bm{\lambda}_{2}\|_{1}.$
By setting $L_{1}=\frac{1}{\rho}\left(L_{r}+\frac{\gamma L_{r}}{1-\gamma}\right)$, we are done. ∎
@@ A.3   Proof of Theorem 2
We prove the Lipschitz continuity of $\Gamma^{n}_{2}$ with respect to both the MF $\mathcal{L}^{n}$ and the policy $\pi^{n}$.
Proof.
Fix $\bm{\lambda}^{n}$, and for notation brevity, we drop the argument of $\bm{\lambda}^{n}$ from the policy $\pi^{n}$. Now we fix $\pi^{n}$ and consider two MF profiles $\mathcal{L}^{n},\mathcal{L}^{n^{\prime}}\in\mathcal{P}(\mathcal{S}\times\mathcal{A})$. Let $\tilde{\Gamma}^{n}_{2}(\mathcal{L}^{n},\pi^{n})(s^{\prime},a^{\prime})=\sum_{s,a}\mathcal{L}^{n}(s,a)\cdot P^{n}(s^{\prime}|s,a^{\prime})\cdot\pi^{n}(a^{\prime}|s).$ We then have,
@@
Now summing over all $(s^{\prime},a^{\prime})\in\mathcal{S}\times\mathcal{A}$:
$\|\tilde{\Gamma}^{n}_{2}(\mathcal{L}^{n},\pi^{n})-\tilde{\Gamma}^{n}_{2}(\mathcal{L}^{n^{\prime}},\pi^{n})\|_{1}=\sum_{s^{\prime},a^{\prime}}\left|\tilde{\Gamma}^{n}_{2}(\mathcal{L}^{n},\pi^{n})(s^{\prime},a^{\prime})-\Gamma^{n}_{2}(\mathcal{L}^{n^{\prime}},\pi^{n})(s^{\prime},a^{\prime})\right|$
@@
$=\sum_{s,a}\mathcal{L}^{n}(s,a)P^{n}(s^{\prime}|s,a^{\prime})\left|\pi^{n}(a^{\prime}|s)-\pi^{n^{\prime}}(a^{\prime}|s)\right|,$
Summing over all $(s^{\prime},a^{\prime})$ gives:
@@
$\leq\sup_{s}\|\pi^{n}(\cdot|s)-\pi^{n^{\prime}}(\cdot|s)\|_{1}\cdot\sum_{s,a}\mathcal{L}^{n}(s,a)=\sup_{s}\|\pi^{n}(\cdot|s)-\pi^{n^{\prime}}(\cdot|s)\|_{1}.$
As a result, we have that
@@
$\leq(1-\zeta)\left(\|\mathcal{L}^{n}-\mathcal{L}^{n^{\prime}}\|_{1}+\sup_{s}\|\pi^{n}(\cdot|s)-\pi^{n^{\prime}}(\cdot|s)\|_{1}\right).$
and we get $L_{2}=L_{3}=1-\zeta$. ∎
@@ A.4   Proof of Theorem 3
Before we present the theorem, we first recall that the definition of contraction mapping and the well-known Banach fixed point theorem state as follows:
@@
Let $(\mathcal{X},d)$ be a non-empty complete metric space. Then a map $T:\mathcal{X}\mapsto\mathcal{X}$ is called a contraction mapping on $\mathcal{X}$ if for any $x,y\in\mathcal{X}$, there exists $c\in[0,1)$ such that $d(T(x),T(y))\leq cd(x,y)$.
@@
Let $(\mathcal{X},d)$ be a non-empty complete metric space with a contraction mapping $T:\mathcal{X}\mapsto\mathcal{X}$. Then $T$ admits a unique fixed point $x^{*}$ in $\mathcal{X}$; that is, $T(x^{*})=x^{*}$.
@@
Proof of Theorem 3. We want to show that the MF forms a contraction mapping. As we are interested in a day-to-day comparison, we would like to check the same entry at two consecutive days at time $t$ and $t+H$ for any $t\geq 0$. With $\mathcal{L}^{n}_{t},\mathcal{L}^{n}_{t+H}\in\mathcal{P}(\mathcal{S}\times\mathcal{A})$, and for each $s^{n}$ of each $n$:
$\|\Gamma^{n}_{2}(\mathcal{L}^{n}_{t},\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t}))-\Gamma^{n}_{2}(\mathcal{L}^{n}_{t+H},\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t+H}))\|_{1}$
$\hskip 5.0pt\leq\|\Gamma^{n}_{2}(\mathcal{L}^{n}_{t},\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t}))-\Gamma^{n}_{2}(\mathcal{L}^{n}_{t},\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t+H}))\|_{1}$
$\hskip 120.0pt+\|\Gamma^{n}_{2}(\mathcal{L}^{n}_{t},\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t+H}))-\Gamma^{n}_{2}(\mathcal{L}^{n}_{t+H},\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t+H}))\|_{1}$
$\hskip 5.0pt\leq L_{3}\|\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t})-\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t+H})\|_{1}+L_{2}\|\mathcal{L}^{n}_{t}-\mathcal{L}^{n}_{t+H}\|_{1}\leq L_{1}L_{3}\|\bm{\lambda}^{n}_{t}-\bm{\lambda}^{n}_{t+H}\|_{1}+L_{2}\|\mathcal{L}^{n}_{t}-\mathcal{L}^{n}_{t+H}\|_{1}$
$\hskip 5.0pt=L_{1}L_{3}|\lambda^{n}_{t}-\lambda^{n}_{t+H}|+L_{2}\|\mathcal{L}^{n}_{t}-\mathcal{L}^{n}_{t+H}\|_{1}\leq(L_{1}L_{\text{MF}}L_{3}+L_{2})\|\mathcal{L}^{n}_{t}-\mathcal{L}^{n}_{t+H}\|_{1}$
where the second last line uses the fact that the only difference between the two vectors occurs at the $T_{\text{hour}}(t)$-th entry. When $L_{1}L_{\text{MF}}L_{3}+L_{2}\in[0,1)$, by the Banach fixed-point theorem, and that $\text{ED}_{\lambda}$ and $\Gamma^{n}_{1}$ are single-valued, we conclude that there exists a unique fixed point of MFE. $\Box$

### v2

@@ 4.1   Components of An Aggregator’s Problem
To promote exploration and improve learning stability, we incorporate entropy regularization into the reward function. Specifically, the single-period reward is modified as:
$r^{\text{REG}}_{n}(s,a,\bm{\lambda}^{n})=r_{n}(s,a,\bm{\lambda}^{n})-\Omega(\pi(\cdot\mid s,\bm{\lambda}^{n})),$ (12)
where $\Omega(\cdot)$ is a $\rho$-strongly convex regularization function. A common choice is the negative entropy: $\Omega(\pi(\cdot\mid s,\bm{\lambda}^{n}))=\alpha\sum_{a\in\mathcal{A}}\pi(a\mid s,\bm{\lambda}^{n})\log\pi(a\mid s,\bm{\lambda}^{n}),$ with $\alpha>0$ controlling the strength of regularization. The corresponding regularized value function is:
$V^{\text{REG}}_{n}(s^{n}_{0},\pi^{n}_{t},\bm{\lambda}^{n})=\mathbb{E}\left[\sum_{\tau=0}^{\infty}\gamma_{n}^{\tau}r^{\text{REG}}_{n}(s^{n}_{\tau},a^{n}_{\tau},\bm{\lambda}^{n})\right].$ (13)
Each aggregator seeks to learn an optimal policy $\pi^{n*}_{t}$ that maximizes $V^{\text{REG}}_{n}$ at each time $t$.
@@ A.1   Uniqueness of Optimal Policy under Regularization
The regularized value function $V^{\text{REG}}_{n}$, as defined in (13), consists of two terms: the expected reward $\mathbb{E}_{a\sim\pi^{n}}[r_{n}(s,a,\bm{\lambda})]$, and a strongly concave regularization term $-\Omega(\pi^{n})$. Since we assume a finite action set, the distribution $\pi^{n}$ corresponds to a probability mass function over a finite set, and the expectation is simply a weighted average of the rewards across actions. Therefore, the expected reward is linear in $\pi^{n}$. Then the sum of a linear function and a strongly concave function remains strongly concave. Hence, the maximizer $\arg\max_{\pi}V^{\text{REG}}_{n}$ is unique. ∎
@@ A.2   Proof of Theorem 1
Proof of Theorem 1.
We consider a family of reinforcement learning problems parameterized by the price vector $\bm{\lambda}$ (aka the LMPs), each corresponding to a distinct $Q$-function. Given $\bm{\lambda}$, we define
$Q_{\bm{\lambda}}(s,a)=r^{\text{REG}}(s,a,\bm{\lambda})+\gamma\sum_{s^{\prime}}Q_{\bm{\lambda}}^{\star}(s^{\prime})\,P(s^{\prime}\mid s,a),$ (22)
where $Q_{\bm{\lambda}}^{\star}(s):=\max_{a}Q_{\bm{\lambda}}(s,a)$ for each state $s\in\mathcal{S}$. Then, starting with the same policy $\pi$, for any $\bm{\lambda}_{1},\bm{\lambda}_{2}$, we have
@@
$=\max_{s,a}\Big{|}r^{\text{REG}}(s,a,\bm{\lambda}_{1})-r^{\text{REG}}(s,a,\bm{\lambda}_{2})+\gamma\sum_{s^{\prime}}Q_{\bm{\lambda}_{1}}^{\star}(s^{\prime})P(s^{\prime}|s,a)-\gamma\sum_{s^{\prime}}Q_{\bm{\lambda}_{2}}^{\star}(s^{\prime})P(s^{\prime}|s,a)\Big{|}$
$\leq\max_{s,a}\Big{\{}|r(s,a,\bm{\lambda}_{1})-r(s,a,\bm{\lambda}_{2})|+\gamma\Big{|}\sum_{s^{\prime}}Q_{\bm{\lambda}_{1}}^{\star}(s^{\prime})P(s^{\prime}|s,a)-\sum_{s^{\prime}}Q_{\bm{\lambda}_{2}}^{\star}(s^{\prime})P(s^{\prime}|s,a)\Big{|}\Big{\}}$
$\leq L_{r}\|\bm{\lambda}_{1}-\bm{\lambda}_{2}\|_{1}+\gamma\|Q_{\bm{\lambda}_{1}}^{\star}-Q_{\bm{\lambda}_{2}}^{\star}\|_{\infty},$
where the first inequality holds by the triangular inequality and the fact that the regularized term $\Omega(\pi)$ cancels out when $\pi$ is the same. For the second inequality, the first term follows the Lipschitz continuity of the reward functions); the second term follows from the fact that $P(\cdot\mid s,a)$ is a probability distribution over $s^{\prime}$, so the absolute difference of weighted sums is bounded by the maximum difference of the weights. By rearranging the terms, we obtain the following result,
$\|Q_{\bm{\lambda}_{1}}^{\star}-Q_{\bm{\lambda}_{2}}^{\star}\|_{\infty}\leq\frac{L_{r}}{1-\gamma}\|\bm{\lambda}_{1}-\bm{\lambda}_{2}\|_{1}$ (23)
Note that by the definition of the regularized reward function in (12), Equation (22) can be rewritten as:
@@
where for any $s$,
$q_{\bm{\lambda}}^{s}:=r(s,\cdot,\bm{\lambda})+\gamma\sum_{s^{\prime}}Q_{\bm{\lambda}}^{\star}(s^{\prime})P(s^{\prime}|s,\cdot)$ (25)
which is $\left(L_{r}+\frac{\gamma L_{r}}{1-\gamma}\right)$-Lipschtiz continuous with respect to $\bm{\lambda}$, as shown below:
@@
$\leq L_{r}\|\bm{\lambda}_{1}-\bm{\lambda}_{2}\|_{1}+\gamma\max_{a}\Big{|}\sum_{s^{\prime}}Q_{\bm{\lambda}_{1}}^{\star}(s^{\prime})P(s^{\prime}|s,a)-\sum_{s^{\prime}}Q_{\bm{\lambda}_{2}}^{\star}(s^{\prime})P(s^{\prime}|s,a)\Big{|}$
$\leq L_{r}\|\bm{\lambda}_{1}-\bm{\lambda}_{2}\|_{1}+\gamma\|Q_{\bm{\lambda}_{1}}^{\star}-Q_{\bm{\lambda}_{2}}^{\star}\|_{\infty}=\left(L_{r}+\frac{\gamma L_{r}}{1-\gamma}\right)\|\bm{\lambda}_{1}-\bm{\lambda}_{2}\|_{1}.$ (26)
Finally, we apply the Fenchel conjugate property stated in Proposition 3. Recall that the optimal policy at state $s$, given belief $\bm{\lambda}$, is defined as $\Gamma_{1}^{n}(s,\bm{\lambda})=\nabla\Omega^{\star}(q_{\bm{\lambda}}^{s}),$ where $\Omega^{\star}$ is the Fenchel conjugate of the regularization function $\Omega$, and $q_{\bm{\lambda}}^{s}$ is the regularized soft Q-function defined in (25). Since $\Omega$ is $\rho$-strongly convex, its conjugate $\Omega^{\star}$ is $\frac{1}{\rho}$-smooth. This means its gradient is Lipschitz continuous with constant $\frac{1}{\rho}$, and thus for any $s\in\mathcal{S}$,
$\|\Gamma^{n}_{1}(s,\bm{\lambda}_{1})-\Gamma^{n}_{1}(s,\bm{\lambda}_{2})\|_{1}\leq\frac{1}{\rho}\|q_{\bm{\lambda}_{1}}^{s}-q_{\bm{\lambda}_{2}}^{s}\|_{\infty}\leq\frac{1}{\rho}\left(L_{r}+\frac{\gamma L_{r}}{1-\gamma}\right)\|\bm{\lambda}_{1}-\bm{\lambda}_{2}\|_{1},$ (27)
where the second inequality comes from (26). Then setting $L_{1}=\frac{1}{\rho}\left(L_{r}+\frac{\gamma L_{r}}{1-\gamma}\right)$ completes the proof. ∎
@@ A.3   Proof of Theorem 2
We prove that the consistency operator $\Gamma^{n}_{2}$, as defined in (14), is Lipschitz continuous with respect to both the MF distribution $\mathcal{L}^{n}$ and the policy $\pi^{n}$.
Proof of Theorem 2.
The consistency operator in (14) consists of two components: (i) the random regeneration of an agent, which occurs with probability $\zeta$, and (ii) the state transition when regeneration does not occur. Since the regeneration step is identical across all agents and independent of the current distribution $\mathcal{L}^{n}$, we focus on the second component – the actual state transition. To simplify notation, we use $\tilde{\Gamma}_{2}^{n}$ to denote only the state transition dynamics:
$\tilde{\Gamma}_{2}^{n}(\mathcal{L}^{n},\pi^{n})(s^{\prime},a^{\prime})=\sum_{s,a}\mathcal{L}^{n}(s,a)\cdot P^{n}(s^{\prime}\mid s,a)\cdot\pi^{n}(a^{\prime}\mid s).$
To proceed with the proof, we fix $\bm{\lambda}^{n}$. For notational convenience, we omit $\bm{\lambda}^{n}$ from the argument of the policy $\pi^{n}$. We first fix $\pi^{n}$ and consider two mean field distributions $\mathcal{L}^{n},\mathcal{L}^{n^{\prime}}\in\mathcal{P}(\mathcal{S}\times\mathcal{A})$. Then we have
@@
Now summing over all $(s^{\prime},a^{\prime})\in\mathcal{S}\times\mathcal{A}$ and using the fact that both $\pi^{n}(\cdot|s)$ and $P^{n}(\cdot|s,a^{\prime})$ are probability mass functions, we have that
$\|\tilde{\Gamma}^{n}_{2}(\mathcal{L}^{n},\pi^{n})-\tilde{\Gamma}^{n}_{2}(\mathcal{L}^{n^{\prime}},\pi^{n})\|_{1}=\sum_{s^{\prime},a^{\prime}}\left|\tilde{\Gamma}^{n}_{2}(\mathcal{L}^{n},\pi^{n})(s^{\prime},a^{\prime})-\tilde{\Gamma}^{n}_{2}(\mathcal{L}^{n^{\prime}},\pi^{n})(s^{\prime},a^{\prime})\right|$
@@
$=\sum_{s,a}\mathcal{L}^{n}(s,a)P^{n}(s^{\prime}|s,a^{\prime})\left|\pi^{n}(a^{\prime}|s)-\pi^{n^{\prime}}(a^{\prime}|s)\right|.$
Summing over all $(s^{\prime},a^{\prime})$ yields
@@
$\leq\sup_{s}\|\pi^{n}(\cdot|s)-\pi^{n^{\prime}}(\cdot|s)\|_{1}\cdot\sum_{s,a}\mathcal{L}^{n}(s,a)=\sup_{s}\|\pi^{n}(\cdot|s)-\pi^{n^{\prime}}(\cdot|s)\|_{1},$
where the last equality holds as the mean-field $\mathcal{L}^{n}(s,a)$, as defined in (9), is a joint probability distribution over $\mathcal{S}\times\mathcal{A}$. Combining the two results above, we have that
@@
$\hskip 10.0pt\leq(1-\zeta)\left(\|\mathcal{L}^{n}-\mathcal{L}^{n^{\prime}}\|_{1}+\sup_{s}\|\pi^{n}(\cdot|s)-\pi^{n^{\prime}}(\cdot|s)\|_{1}\right),$
and we let $L_{2}=L_{3}=1-\zeta$. ∎
@@ A.4   Proof of Theorem 3
Before presenting the theorem, we first recall the definition of a contraction mapping and the classical Banach fixed-point theorem:
@@
Let $(\mathcal{X},d)$ be a non-empty complete metric space. A map $T:\mathcal{X}\mapsto\mathcal{X}$ is a contraction mapping on $\mathcal{X}$ if $\;\forall x,y\in\mathcal{X}$, there exists $c\in[0,1)$ such that $d(T(x),T(y))\leq cd(x,y)$.
Then, we present the Banach fixed point theorem:
@@
Let $(\mathcal{X},d)$ be a non-empty complete metric space with a contraction mapping $T:\mathcal{X}\mapsto\mathcal{X}$. Then $T$ admits a unique fixed point $x^{*}$ in $\mathcal{X}$. That is, $T(x^{*})=x^{*}$.
@@
Proof of Theorem 3. The key here is to show that the consistency operator $\Gamma^{n}_{2}$ is a contraction mapping. To do so, we focus on comparing the distributions at the same time of day across consecutive days. Let $H$ denote the length of a full day, and fix any time index $t\geq 0$. Due to the presence of diurnal patterns, it is natural to analyze the evolution of the mean-field distributions at times $t$ and $t+H$, corresponding to the same time period on two successive days. Let $\mathcal{L}^{n}_{t},\mathcal{L}^{n}_{t+H}\in\mathcal{P}(\mathcal{S}\times\mathcal{A})$ denote the mean-field distributions for aggregator $n$ at these two time points. We aim to establish that the update operator is contractive with respect to the $\ell_{1}$-norm over the joint state–action distribution:
$\|\Gamma^{n}_{2}(\mathcal{L}^{n}_{t},\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t}))-\Gamma^{n}_{2}(\mathcal{L}^{n}_{t+H},\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t+H}))\|_{1}{\leq}\|\Gamma^{n}_{2}(\mathcal{L}^{n}_{t},\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t}))-\Gamma^{n}_{2}(\mathcal{L}^{n}_{t},\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t+H}))\|_{1}$
$\hskip 200.0pt+\|\Gamma^{n}_{2}(\mathcal{L}^{n}_{t},\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t+H}))-\Gamma^{n}_{2}(\mathcal{L}^{n}_{t+H},\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t+H}))\|_{1}$
$\hskip 5.0pt\overset{(a)}{\leq}L_{3}\|\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t})-\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t+H})\|_{1}+L_{2}\|\mathcal{L}^{n}_{t}-\mathcal{L}^{n}_{t+H}\|_{1}\overset{(b)}{\leq}L_{1}L_{3}\|\bm{\lambda}^{n}_{t}-\bm{\lambda}^{n}_{t+H}\|_{1}+L_{2}\|\mathcal{L}^{n}_{t}-\mathcal{L}^{n}_{t+H}\|_{1}$
=$\hskip 5.0pt\overset{(c)}{=}L_{1}L_{3}|\lambda^{n}_{t}-\lambda^{n}_{t+H}|+L_{2}\|\mathcal{L}^{n}_{t}-\mathcal{L}^{n}_{t+H}\|_{1}\overset{(d)}{\leq}(L_{1}L_{\text{MF}}L_{3}+L_{2})\|\mathcal{L}^{n}_{t}-\mathcal{L}^{n}_{t+H}\|_{1}.$
In the above derivations, inequality (a) follows from Theorem 2; while inequality (b) follows from Theorem 1. For the equality (c), it uses the fact that the only difference between the two vectors occurs at the $T_{\text{hour}}(t)$-th entry. The last inequality follows from the Lipschitz continuity of the LMPs as stated in Proposition 1. If the constant $L_{1}L_{\text{MF}}L_{3}+L_{2}\in[0,1)$, then the composition of the update operators defines a contraction mapping under the $\ell_{1}$-norm. Since $\lambda^{n}_{t}(\cdot)$ and the optimality operator $\Gamma^{n}_{1}$ are both single-valued, Banach’s fixed-point theorem guarantees the existence and uniqueness of the mean field equilibrium. $\Box$